# Anatomy of spin-orbit-torque-assisted magnetization dynamics in Co/Pt bilayers: Importance of the orbital torque

### v1

@@ II.5 Atomistic spin dynamics simulations
where $\gamma=1.76\times 10^{11}\leavevmode\nobreak\ \textrm{s}^{-1}\textrm{T}^{-1}$ is the gyromagnetic ratio and $\alpha$ the dimensionless damping constant. Temperature is included via Langevin dynamics by adding a Gaussian white noise $\bm{\xi}_{i}$ to the effective field $\bm{H}_{i}^{\rm eff}=-{\frac{\partial\mathcal{H}}{\partial{\bm{S}_{i}}}}+\bm{\xi}_{i}$. The two terms in the SLLG equation describe precession and damping dynamics through the torques acting on the spin system.
@@
where $\bm{H}^{\textrm{SOT}}_{i}$ is the field coming from the interaction of the strong moments with electrically-induced moments. In our model, these interactions are incorporated in the extended Heisenberg Hamiltonian, Eq. (13), with $-\partial\mathcal{H}/\partial\bm{S}_{i}=\bm{H}^{\textrm{eff}}_{i}+\bm{H^{\textrm{SOT}}}_{i}$, where the first and third term in the Hamiltonian result in the effective field $\bm{H}^{\textrm{eff}}_{i}$ from strong moments and the remaining terms collectively give $\bm{H^{\textrm{SOT}}}_{i}$ on differentiating the Hamiltonian with respect to the strong moments. From Eq. (13), it can be clearly seen that $\bm{H^{\textrm{SOT}}}_{i}$ is a linear function of the induced moments and $\bm{T}^{\textrm{SOT}}_{i}$ can, hence, be represented as
@@
$\bm{T^{\textrm{SOT}}}_{i}$ $\propto P(A)\,\bm{S}_{i}\times(\bm{E}\times\hat{\bm{z}})$
$-P(B)\,\bm{S}_{i}\times(\bm{S}\times(\bm{E}\times\hat{\bm{z}}))$
$+P(A^{\prime})\,(\bm{S}\cdot\bm{E})\bm{S}_{i}\times(\bm{S}\times\hat{\bm{z}})$
@@ III.1 Switching behavior
According to our ab initio calculations our model shows an easy-plane anisotropy. Therefore, the initial magnetization direction in the cobalt is set to the in-plane $\hat{\bm{x}}$-direction. We compare the magnetization dynamics in the presence or absence of induced orbital moments in addition to the induced spin moments for varying electric field magnitudes. As is evident from Eq. (8), the induced moments and with that the SOTs are directly proportional to the electric field amplitude.
@@
To start with, we study the dynamics of our model in the presence of all the induced moments and compare with the absence of induced orbital moments. As can be seen in Fig. 3, in the presence of induced orbital moments the $90$-degree in-plane switching begins to appear at comparatively lower electric field magnitude ($E=0.8$ mV/nm), where the switching direction is given by $\bm{E}\times\hat{\bm{z}}$. The SOT switches the magnetization from the $\hat{\bm{x}}$-direction to the $-\hat{\bm{y}}$-direction. As the switching direction is in-plane and since the system has easy-plane anisotropy only, i. e., no anisotropy energy barrier in plane, switching occurs when SOT effects overcome the thermal fluctuations. The influence of the induced spin moments alone ($\delta\mu_{L}=0$) is much weaker as now the switching can be observed only for higher electric fields ($E\geq 2$ mV/nm).
@@
On further increasing the electric-field strength it can be observed that the system first exhibits faster switching ($E=2$ mV/nm and $10$ mV/nm) and then begins to oscillate at even higher fields. To understand this peculiar behavior, we have studied the switching behavior under the influence of M-even and M-odd dependent induced moments separately.
As expressed in Eqs. (14) and (15), the induced moments have magnetization-independent ($P(A)$ term) and magnetization-dependent $M$-even (terms with $P(A^{\prime})$, $P(A^{\prime\prime})$ and $P(A^{\prime\prime\prime})$) and $M$-odd (terms with $P(B)$ and $P(B^{\prime})$) origins. The influence of $M$-even and $M$-odd induced moments can nonetheless be separately studied.
As depicted in Fig. 4, it is the $M$-even induced moments that support switching in the $\bm{E}\times\hat{\bm{z}}$ direction at any electric field magnitude, whereas the $M$-odd moments are weaker in magnitude and begin to contribute significantly only at higher electric fields (here, $E=20$ mV/nm). Moreover, these $M$-odd moments support a switching direction given by $-\bm{S}\times(\bm{E}\times\hat{\bm{z}})$ as well as $(\bm{S}\cdot\bm{E})\bm{\hat{z}}$, which thus applies torque in the out-of-plane direction, facing limits set by the anisotropy of the system and applies an anti-damping torque that moves the magnetization towards $+\hat{\bm{y}}$ direction, if stronger than the thermal energy (at higher field $E=20$ mV/nm), as we will see in the torque section in detail.
@@
The time scale of the switching process can be quantified as the time it takes for the magnetization to reach $-\bm{\hat{y}}$ with $M_{x}=0$ for the first time after having applied the field. For the cases discussed above, it varies from $90$ ps (at low electric field) to $13$ ps (at $E=10$ mV/nm) for magnetization-dependent moments when the induced orbital moments are present, whereas slower rates are obtained in the absence of induced orbital moments, as can be expected (see Sec. V in Supplemental Material [53]).
@@
In the case of magnetization-independent moments, the switching time is almost the same
with or without induced orbital moments and faster than for the magnetization dependent case. Moreover, switching occurs with only magnetization-independent spin moments for lower electric fields than in the presence of magnetization-dependent spin moments where it does not switch (see Sec. V in Supplemental Material [53]). This means that the magnetization-dependent induced moments are crucial to explain the effect of induced orbital moments on the switching, as will be discussed in the next subsection in connection with the layer-resolved torques.
@@ III.2 Layer-resolved spin and orbital torque
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\alpha\bm{S}_{i}\times\left(\bm{S}_{i}\times\left(T_{\text{F}}\left(\bm{P}\right)+T_{\text{D}}\left(-\bm{S}\times\bm{P}\right)\right)\right).$ (25)
@@
an initial damping-like torque acting in direction of $\bm{P}$, is referred to as a damping torque whereas the torque acting in $-\bm{P}$ direction is referred to as being anti-damping.
@@
As shown in Fig. 6 (right), in this case the field-like torque has strong contributions from the orbital torque arising from both, the interfacial Pt and Co layers, thus enhancing the overall field-like torque compared to the case which induces the effects of the Co/vacuum surface layer. This leads to faster switching with a switching time of $45$ ps, as can be seen from magnetization dynamics, whereas switching observed in the presence of Co/vacuum surface moments takes $90$ ps. This furthermore indicates that in absence of current-induced Co/vacuum surface moments the switching can occur at even lower electric field strengths. Specifically, we have observed switching at a minimum field of $E_{x}=0.4$ mV/nm with a switching time of $85$ ps, described in detail in Sec. VII of the Supplemental Material [53]. Lastly, we have also analyzed the SOT-field dynamics for the SOT-interaction terms in Eq. (13) and observed that the SOT field from the $M$-even induced spin moments is strongly magnetization dependent whereas the SOT field from $M$-even orbital moment interaction has negligibly small variations with changes of the in-plane magnetization directions, see Sec. VIII of the Supplemental Material [53]. This explains why our simulations uncover a large role played by the $M$-even
@@ IV Discussion and Conclusions
The spin part of the SOT has already been widely investigated (see [43]), yet not in a layer-resolved manner. Also, not much is known about the orbital contribution to the SOT and how it operates. Orbital torques have been proposed in NM/FM bilayers [45, 70, 71, 72, 73], but in experiments, such as spin-torque ferromagnetic resonance measurements, it is often difficult to separate spin and orbital torque contributions. Recently, Wang et al. [70] deduced a large damping-like torque due to the OREE in Pt/CoO bilayers, and Krishnia et al. [72] found a large damping-like torque in the Pt/Co/Cu/CuO system due to the OREE. In a different heterostructure, Pt/Co/Al/Pt, Krishnia et al. [74] found a large enhancement of the field-like SOT due to orbital texture and OREE at the Co/Al interface [75]. Consistent with this, we find in our study a large field-like torque due to accumulation of orbital moments at the Pt/Co interface, even though in general, the SOT appears to depend on the specifics of the heterostructure. Our simulated results are furthermore in agreement with a recent experimental study of switching in Pt/CoFeB by Zhang et al. [76], who found that field-like SOT originating from the ferromagnetic layer can dominate the switching process. The thus-induced switching could be fast ($\sim 0.2$ ns) and occur at relatively low electric field strengths [76], consistent with our findings.
@@
Since our analysis provides layer-resolved spin-orbit torques, it can be concluded that the torque applied through induced spin moments at the interface layer by exchange interaction has a major contribution to an anti-damping behavior, arising from the $M$-odd induced spin moments. The in-plane switching occurs due to the action of a field-like SOT applied primarily via the $M$-even induced orbital moments at the Pt/Co interface. These are of nonrelativistic origin, but lead to an orbital torque due to the spin-orbit interaction. The spin-orbit interaction from the Pt interface layer is here taken into account via renormalization and is confirmed to be an important term supporting switching.
@@
by a suitable interface engineering. Moreover, we have observed that on turning off the current-induced moments at the Co/vacuum surface, the switching occurs on a faster time-scale, ($45$ ps at $E_{x}=0.8$ mV/nm) and can occur at even lower electric fields of magnitude $E_{x}=0.4$ mV/nm with switching time of $85$ ps. Finally, we conclude that it is the large $M$-even
@@ Acknowledgements.
We thank Daegeun Jo for valuable discussions. This work has been supported by the the German Research Foundation (Deutsche Forschungsgemeinschaft) through CRC/TRR 227 (project MF, project-ID: 328545488)) and CRC 1432 (project B02, project-ID: 425217212), the Swedish Research Council (VR), the K. and A. Wallenberg Foundation (Grants No. 2022.0079 and 2023.0336), by the EIC Pathfinder OPEN grant No. 101129641 “OBELIX”, by the National Research, Development, and Innovation Office (NRDI) of Hungary under Project Nos. K131938, K142652, and FK142601, by the Thematic Area Excellence Program of the Ministry of Culture and Innovation from the NRDI Fund through the Grant No. TKP2021-NVA-02, and by the Hungarian Academy of Sciences via a János Bolyai Research Grant (Grant No. BO/00178/23/11). Computational resources were provided by the core facility SCCKN in Konstanz and the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at NSC Linköping, partially funded by VR through Grant Agreement No. 2022-06725.

### v2

@@ II.5 Atomistic spin dynamics simulations
where $\gamma=1.76\times 10^{11}\penalty 10000\ \textrm{s}^{-1}\textrm{T}^{-1}$ is the gyromagnetic ratio and $\alpha$ the dimensionless damping constant. Temperature is included via Langevin dynamics by adding a Gaussian white noise $\bm{\xi}_{i}$ to the effective field $\bm{H}_{i}^{\rm eff}=-{\frac{\partial\mathcal{H}}{\partial{\bm{S}_{i}}}}+\bm{\xi}_{i}$. The two terms in the SLLG equation describe precession and damping dynamics through the torques acting on the spin system.
@@
where $\bm{H}^{\textrm{SOT}}_{i}$ is the field coming from the interaction of the strong moments with electrically-induced moments. In our model, these interactions are incorporated in the extended Heisenberg Hamiltonian, Eq. (13), with $\partial\mathcal{H}/\partial\bm{S}_{i}=\bm{H}^{\textrm{eff}}_{i}+\bm{H^{\textrm{SOT}}}_{i}$, where the first and third term in the Hamiltonian result in the effective field $\bm{H}^{\textrm{eff}}_{i}$ from strong moments and the remaining terms collectively give $\bm{H^{\textrm{SOT}}}_{i}$ on differentiating the Hamiltonian with respect to the strong moments. From Eq. (13), it can be clearly seen that $\bm{H^{\textrm{SOT}}}_{i}$ is a linear function of the induced moments and $\bm{T}^{\textrm{SOT}}_{i}$ can, hence, be represented as
@@
[…]
$)\displaystyle-P(B)\,\bm{S}_{i}\times(\bm{S}\times(\bm{E}\times\hat{\bm{z}}$))
[…]
@@ III.1 Switching behavior
According to our ab initio calculations our model shows an easy-plane anisotropy. Therefore, the initial magnetization direction in the cobalt is set to the in-plane $\hat{\bm{x}}-$direction. We compare the magnetization dynamics
in the presence or absence of induced orbital moments in addition to the induced spin moments for varying electric field magnitudes. As is evident
from Eq. (8), the induced moments and with that the SOTs are directly proportional to the electric field amplitude.
@@
To start with, we study the dynamics of our model in the presence of all the induced moments and compare with the absence of induced orbital moments. As can be seen in Fig. 3, in the presence of induced orbital moments ($\delta\mu_{L}\neq 0$) the $90$-degree in-plane switching begins to appear at comparatively lower electric field magnitude ($E=0.8$ mV/nm), where the switching direction is given by $\bm{E}\times\hat{\bm{z}}$. The SOT switches the magnetization from the $\hat{\bm{x}}-$direction to the $\hat{\bm{y}}-$direction. As the switching direction is in-plane and since the system has easy-plane anisotropy only, i. e., no anisotropy energy barrier in plane, switching occurs when SOT effects overcome the thermal fluctuations, There is, however, no final steady state as long as the SOT induced field is weaker than the thermal fluctuations and the magnetization vector will fluctuate in the easy plane as can be seen at low electric field ($E=0.8$ mV/nm) and in absence of induced orbital moments ($\delta\mu_{L}=0$). The influence of the induced spin moments alone ($\delta\mu_{L}=0$) is much weaker as now the switching can be observed only for higher electric fields ($E\geq 2$ mV/nm).
@@
On further increasing the electric-field strength it can be observed that the system first exhibits faster switching ($E=2$ mV/nm and $10$ mV/nm), and then begins to oscillate at even higher fields again not reaching a final steady state. Finally, for fields of $E=2$ mV/nm and $10$ mV/nm the system switches in $\hat{\bm{y}}$ direction for both, presence and absence of induced orbital moments, but with differences in switching and relaxation times. E.g., for $E=10$ mV/nm the magnetization has reached its final state only in the presence of induced orbital moments, while the magnetization is still precessing on the same time scale in the absence of induced orbital moments.
To understand the peculiar oscillating behavior for higher fields, we have separated the influence of $M$-even and $M$-odd dependent induced moments. As expressed in Eqs. (14) and (15), the induced moments have magnetization-independent ($P(A)$ term) and magnetization-dependent $M$-even (terms with $P(A^{\prime})$, $P(A^{\prime\prime})$ and $P(A^{\prime\prime\prime})$) and $M$-odd (terms with $P(B)$ and $P(B^{\prime})$) origins. The influence of $M$-even and $M$-odd induced moments can nonetheless be separately studied.
[…]
@@
The time scale of the switching process can be quantified as the time it takes for the magnetization to reach $\bm{\hat{y}}$ with $M_{x}=0$ for the first time after having applied the field. For the cases discussed above, it varies from $90$ ps (at low electric field) to $13$ ps (at $E=10$ mV/nm) for magnetization-dependent moments when the induced orbital moments are present, whereas slower rates are obtained in the absence of induced orbital moments, as can be expected (see Sec. V in Supplemental Material [53]).
@@
In the case of magnetization-independent moments, the switching time is almost the same with or without induced orbital moments and faster than for the magnetization dependent case. Moreover, switching occurs with only magnetization-independent spin moments for lower electric fields than in the presence of magnetization-dependent spin moments where it does not switch (see Sec. V in Supplemental Material [53]). This means that the magnetization-dependent induced moments are crucial to explain the effect of induced orbital moments on the switching, as will be discussed in the next subsection in connection with the layer-resolved torques.
@@ III.2 Layer-resolved spin and orbital torque
$\penalty 10000\ \penalty 10000\ \penalty 10000\ -\alpha\bm{S}_{i}\times\left(\bm{S}_{i}\times\left(T_{\text{F}}\left(\bm{P}\right)+T_{\text{D}}\left(-\bm{S}\times\bm{P}\right)\right)\right).$ (25)
@@
an initial damping-like torque acting in direction of $\bm{P}$, is referred to as a damping torque whereas the torque acting in $\bm{P}$ direction is referred to as being anti-damping.
@@
As shown in Fig. 6 (right), in this case the field-like torque has strong contributions from the orbital torque arising from both, the interfacial Pt and Co layers, thus enhancing the overall field-like torque compared to the case which induces the effects of the Co/vacuum surface layer. This
leads to faster switching with a switching time of $45$ ps, as can be seen from magnetization dynamics, whereas switching observed in the presence of Co/vacuum surface moments takes $90$ ps. This furthermore indicates that in absence of current-induced Co/vacuum surface moments the switching can occur at even lower electric field strengths. Specifically, we have observed switching at a minimum field of $E_{x}=0.4$ mV/nm with a switching time of $85$ ps, described in detail in Sec. VII of the Supplemental Material [53]. Lastly, we have also analyzed the SOT-field dynamics for the SOT-interaction terms in Eq. (13)
and observed that the SOT field from the $M$-even induced spin moments is strongly magnetization dependent whereas the SOT field from $M$-even orbital moment interaction has negligibly small variations with changes of the in-plane magnetization directions, see Sec. VIII of the Supplemental Material [53]. This explains why our simulations uncover a large role played by the $M$-even
@@ IV Discussion and Conclusions
The spin part of the SOT has already been widely investigated (see [43]) often in a macrospin approximation, yet not in a layer-resolved manner. Also, not much is known about the orbital contribution to the SOT and how it operates. Hence, our atomic-layer and spin- and orbital-resolved approach provides additional insight. Orbital torques have been proposed in NM/FM bilayers [45, 70, 71, 72, 73], but in experiments, such as spin-torque ferromagnetic resonance measurements, it is often difficult to separate spin and orbital torque contributions. Recently, Wang et al. [70] deduced a large damping-like torque due to the OREE in Pt/CoO bilayers, and Krishnia et al. [72] found a large damping-like torque in the Pt/Co/Cu/CuO system due to the OREE. In a different heterostructure, Pt/Co/Al/Pt, Krishnia et al. [74] found a large enhancement of the field-like SOT due to orbital texture and OREE at the Co/Al interface [75]. Consistent with this, we find in our study a large field-like torque due to accumulation of orbital moments at the Pt/Co interface, even though in general, the SOT appears to depend on the specifics of the heterostructure. Our simulated results are furthermore in agreement with a recent experimental study of switching in Pt/CoFeB by Zhang et al. [76], who found that field-like SOT originating from the ferromagnetic layer can dominate the switching process. The thus-induced switching could be fast ($\sim 0.2$ ns) and occur at relatively low electric field strengths [76], consistent with our findings.
@@
Since our analysis provides layer-resolved spin-orbit torques, it can be concluded that the torque applied through induced spin moments at the interface layer by exchange interaction has a major contribution to an anti-damping behavior, arising from the $M$-odd induced spin moments. The in-plane switching occurs due to the action of a field-like SOT applied primarily via the $M$-even induced orbital moments at the Pt/Co interface. From earlier work on current-induced orbital moments in metallic bilayers [27], it is known that these $M$-even induced orbital moments are independent of the strength of spin-orbit interaction and are thus of nonrelativistic origin. However, they can apply a torque on the local spin moments only through the local spin-orbit interaction. Thus, although being of nonrelativistic nature the induced orbital moments require spin-orbit interaction with local spin moments to apply a torque. The spin-orbit interaction from the Pt interface layer is taken into account via renormalization and is confirmed to be an important term supporting switching.
We note that the net field-like torque for Pt/Co is larger than the (anti-)damping-like torque, in contrast to most previous Pt/Co experimental studies that consider comparatively thicker Pt layer. From our first-principles calculations it is found that the $M$-even induced spin moments (pointing in the $\bm{E}\times\hat{\bm{z}}$ direction) have contribution arising mostly from the interface, which in turn applies a field-like torque, and the $M$-odd induced spin moments (pointing in $\bm{M}\times\bm{E}\times\hat{\bm{z}}$ direction) have a mixed bulk and interface origin and apply an anti-damping like torque. Indeed, for the spin part, due to filtering and scattering effects at the interface, the net field-like contribution at the Co layers diminishes and the anti-damping torque from the spin part gets stronger. This is usually the case for thick Pt layers where the bulk SHE contribution dominates and generates stronger damping-like torque. In experiments [77], it has indeed been shown that on decreasing the Pt thickness the bulk SHE contribution decreases leading to a reduced in damping-like effective-field. Moreover, it is difficult to clearly separate the effective field-like and damping-like fields for smaller Pt thickness where interface effects begin to dominate and a net increment in field-like torque begins. In our calculations it has also been shown that the strong orbital contribution comes from the interface effects which again explains that strong-field-like torque clearly dominates in a thin Pt/Co bilayer. Previously, it has been found [45] that the dominating torques from induced orbital moments and induced spin moments have the same direction. In our case, these similar directions can be explained from the direction of the net effective field-like torque from $M$-even induced spin moments at the Co and Pt interface layer and $M$-even induced orbital moments, leading to switching in the same direction.
Additionally, it should be noted that the effect of spin diffusion is already present in the magnetoelectric susceptibility tensor computed via relativistic first-principles approach. However, it might indeed be a question whether the influence of spin diffusion would not be more pronounced when one considers a thick Pt layer, which in turn might enhance the effect of damping-like torque.
@@
by a suitable interface engineering. Moreover, we have observed that on turning off the current-induced moments at the Co/vacuum surface, the switching occurs on a faster time-scale, ($45$ ps at $E_{x}=0.8$ mV/nm) and can occur at even lower electric fields of magnitude $E_{x}=0.4$ mV/nm with switching time of $85$ ps. Finally,
we conclude that it is the large $M$-even
@@ Acknowledgements.
We thank Daegeun Jo for valuable discussions. This work has been supported by the the German Research Foundation (Deutsche Forschungsgemeinschaft) through CRC/TRR 227 (project MF, project-ID: 328545488)) and CRC 1432 (project B02, project-ID: 425217212), the Swedish Research Council (VR), the K. and A. Wallenberg Foundation (Grants No. 2022.0079 and 2023.0336), by the EIC Pathfinder OPEN grant No. 101129641 “OBELIX”, by the National Research, Development, and Innovation Office (NRDI) of Hungary under Project Nos. K131938, K142652, FK142601 and ADVANCED 149745, by the Thematic Area Excellence Program of the Ministry of Culture and Innovation from the NRDI Fund through the Grant No. TKP2021-NVA-02, and by the Hungarian Academy of Sciences via a János Bolyai Research Grant (Grant No. BO/00178/23/11). Computational resources were provided by the core facility SCCKN in Konstanz and the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at NSC Linköping, partially funded by VR through Grant Agreement No. 2022-06725.